OpenStudy (anonymous):

-4y + 17 <= 29 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (anonymous):

thats part one. thank you! can you please do part 2

OpenStudy (tkhunny):

That is not Part 1: After obtaining \(-4y \le 12\), solve two ways, just to prove that it is as always stated. Add 4y \(0 \le 12+4y\) Subtract 12 \(-12 \le 4y\) Divide by 4 \(-3 \le y\) Contract that with: \(-4y \le 12\) Divide by -4 and since that is negative, reverse the inequality. \(y \ge -3\)

OpenStudy (anonymous):

 thats how i do it